Depth of Bore
Edge condition Above target Below target
coarse 15 10
moderate 21 24
smooth 49 81

The data from 200 machined parts are summarized as follows ^

What is the probability that a part selected does not have a moderate edge condition or does not have a below-target bore depth? Round your answer to two decimal places.
Answer #1

P(doesn't have moderate edge condition or doesn't have a below-target bore depth) = P(doesn't have moderate edge condition) + P(doesn't have a below-target bore depth) - P(doesn't not have moderate edge condition and doesn't have below-target bore depth)

= (15 + 49 + 10 + 81)/200 + (15 + 21 + 49)/200 - (15 + 49)/200

= 155/200 + 85/200 - 65/200

= 175/200 = 0.88